What are the asymptotes of the hyperbola given by the equation (y^2/1)-(x^2/121)=1?
Answer by lwsshak3(11628)

What are the asymptotes of the hyperbola given by the equation
(y^2/1)-(x^2/121)=1?
**
This is an equation of a hyperbola with vertical transverse axis.
Its standard form: (y-k)^2/a^2-(x-h)^2)/b^2=1, (h,k)=(x,y) coordinates of center
For given equation: (y^2/1)-(x^2/121)=1
center: (0,0)
a^2=1
a=1
b^2=121
b=√121=11
..
Asymptotes are straight lines that go thru the center (0,0)
Standard form of equation for straight lines: y=mx+b, m=slope, b=y-intercept
For hyperbolas with vertical transverse axis:
slopes of asymptotes=±a/b=1/11
equations of asymptotes:
y=±x/11+b
since center is at (0,0), y-intercept, b=0
equation of asymptotes:
y=x/11
and
y=-x/11